Suppose Johnson & Johnson and the Walgreen Company have the expected returns and volatilities shown below, with a correlation of 21%.

E[R] SD[R]
Johnson & Johnson 5% 14%
Walgreen Company 10% 20%

Consider a portfolio that is equally invested in Johnson & Johnson's and Walgreen's stock

(a) Calculate the expected return as a percent.

%

(b) Calculate the volatility (standard deviation) of returns as a percent. (Round your answer to two decimal places.)

%
Answer #1

Given

Expected return of Johnson & Johnson E[r1]= 5%

Expected return of Walgreen Company E[r2]= 10%

Standard Deviation of Johnson & Johnson SD[r1]= 14%

Standard Deviation of Walgreen Company SD[r2]= 10%

Correlation between 2 stock [corr] = 21%

For a portfolio with equally invested in Johnson & Johnson's and Walgreen's stock, Expected return is weighted average

a). E[rp] = w1*E[r1] + w2*E[r2] = 0.5*5 + 0.5*10 = 7.5%

b). Standard deviation can be calculated using following formula,

SD[rp] = SQRT((SD[r1]^2) + (SD[r2]^2) + 2*w1*w2*corr*SD[r1]*SD[r2]) = (14^2 + 10^2 + 2*0.5*0.5*.21*14*10)^(1/2) = 17.63%